# Reflectance Prediction-based Knowledge Distillation for Robust 3D Object Detection in Compressed Point Clouds

### v1

@@ IV-B Experimental Results
Table I presents the performance improvement of our RPKD framework in 3D object detection for compressed point clouds with the PV-RCNN and Voxel-RCNN backbones on the KITTI validation split. For the compressed point clouds of PCC-S-C12, PCC-S-C11, and PCC-S-C10, RPKD-PV increases mAP by 3.33, 4.23, and 3.65, while RPKD-Voxel improves mAP by 2.25, 3.22, and 3.3. As a result, it reveals that the detectors of our RPKD framework exhibit consistent effectiveness and robustness across multiple code rates and backbones. Additionally, for compressed point clouds with the same code-rate order, RPKD-PV boosts the AP values for moderate-level pedestrians by 9.02, 11.98, and 9.66, respectively, and RPKD-Voxel elevates the AP values for moderate-level pedestrians by 7.93, 8.59, and 7.36. These findings highlight the significant enhancement in 3D detector robustness for challenging small-sized pedestrians within compressed point clouds.
@@
As depicted in Fig. 5(a), based on the PV-RCNN backbone, we compare the detection performance of PV-RCNN (baseline), existing SparseKD, simulated RPV-KD and our RPKD on the KITTI compressed point clouds across multiple code rates. RPKD-PV achieves higher mAP values in both medium and high code rates, and its mAP value in the medium code rate exceeds that of RP-PV (T2). Fig. 5(b) shows the generalization performance of our RPKD framework on the Voxel-RCNN backbone. RPKD-Voxel outperforms Voxel-RCNN (baseline) in all code rates, and its mAP value surpasses that of RP-Voxel (T3) and Voxel-RCNN backbone in the high code rate. These results vividly illustrate the outstanding robustness and remarkable enhancement capabilities of our method in 3D object detection for compressed point clouds.
For the generalization ability in WOD, we evaluate the detection performance of our RPKD framework across multiple code rates with the PV-RCNN backbone on the validation split. In Table II, RPKD-PV improves mAP/mAPH for PCC-S-C12, PCC-S-C11, and PCC-S-C10 by 0.29/2.1, 0.26/1.77, and 0.86/1.48, respectively. For the compressed point clouds of PCC-S-C12, the APH values for vehicle and pedestrian categories exceed those of RP-PV (T2) and PV-RCNN backbone on raw data. The experimental results demonstrate that our RPKD framework maintains consistent effectiveness throughout diverse code rates, backbone networks and detection datasets.
@@
| Voxel-wise | Point-wise | PV-RCNN | RP-PV | |||
@@ IV-C1 Different Configurations of RP Module
| Data Type | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | |||
| Raw Data | PV-RCNN | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.56 |
| RP-PV(T2) | 92.10 | 84.36 | 82.29 | 63.62 | 56.35 | 50.88 | 90.30 | 70.75 | 66.96 | 73.07 | |
| PCC-S-C12 (Bpp=3.864) | PV-RCNN(Baseline) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
| RPKD-PV(Unified) | 91.46 | 82.92 | 80.01 | 65.55 | 58.14 | 53.16 | 90.54 | 67.30 | 62.98 | 72.45 | |
| PCC-S-C11 (Bpp=2.146) | PV-RCNN(Baseline) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
| RPKD-PV(Unified) | 91.20 | 82.37 | 79.29 | 66.71 | 59.32 | 54.18 | 87.66 | 66.86 | 62.26 | 72.21 | |
| PCC-S-C10 (Bpp=1.047) | PV-RCNN(Baseline) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
| RPKD-PV(Unified) | 91.37 | 82.37 | 79.48 | 65.38 | 58.05 | 52.91 | 85.69 | 61.66 | 57.44 | 70.48 | |
@@ IV-C2 Critical Elements of Our RPKD
As indicated in Table IV, we assessed the influences of critical elements in our RPKD framework on the detection performance for compressed point clouds. In the first two rows, we compared the pre-training performance of compressed point-cloud detectors, where refers to the voxel-based averaging operation in the RCM module. denotes screening foreground compressed points in calculating reflectance distillation loss. In the third row, we used the matched raw point reflectance as the distillation label during point-wise RKD and did not perform the voxel-based . In contrast, we utilized the matched reflectance of raw point-cloud voxels as the distillation label during voxel-wise RKD and conducted the voxel-based , as shown in the fourth row. In Table IV, PV-RCNN and RP-PV represent raw point-cloud detector types that provided the first-stage proposals for DKD. As indicated in the last row, our RPKD framework achieves the best performance with these settings.
@@ IV-C3 Designed Modules of Our RPKD
As shown in Table V, we evaluated the designed module’s effectiveness in 3D object detection for compressed point clouds. The first row displays the detection performance of PV-RCNN (baseline) for compressed point clouds of PCC-S-C11. We then incrementally introduced the RP module, RKD and DKD into the detector and reported corresponding mAP values. All the designed modules improve the performance of PV-RCNN (baseline), with the best mAP value achieved when all modules are utilized together.
IV-C4 Experiments for a Unified Model
To develop a unified detector applicable to various code rates, we combined the KITTI training splits of PCC-S-C10, PCC-S-C11, and PCC-S-C12 in a ratio as input for the unified model. The trained model was then used to perform inference on each validation split of these compressed point clouds to evaluate detection performance. As shown in Table VI, RPKD-PV (unified) increases mAP by 2.84 for both PCC-S-C12 and PCC-S-C11, and by 4.04 for PCC-S-C10. These results demonstrate that our RPKD framework under a unified model achieves consistent effectiveness across diverse code rates.
IV-C5 Visual Comparisons of Detection Results
As depicted in Fig. 6, we performed visual comparisons of different data types and detection methods on the KITTI validation split. For compressed point-cloud detectors, RPKD-PV achieves fewer false positives and missed detections across multiple code rates, demonstrating the stronger robustness of our RPKD framework in 3D object detection for compressed point clouds. However, compared to the raw point-cloud detector based on PV-RCNN, RPKD-PV tends to miss objects with fewer points or at farther distances, which reveals a limitation of our method in detecting such challenging objects.
IV-C6 Visual Comparisons of Reflectance Prediction
In each BEV grid, we selected the maximum reflectance value from either raw data or compressed point clouds for coloring the corresponding visual points. As shown in Fig. 7, the predicted reflectance in compressed point clouds closely matches the actual reflectance in raw data within the major reflective areas of foreground objects. These results demonstrate that our RP module accurately predicts the reflectance of foreground objects across various code rates, thereby improving detection robustness for compressed point clouds.
IV-D Discussion of Limitation and Future Work
For non-reflectance compressed point clouds, this paper effectively supplements the missing reflectance information through the RP module and RKD, thus enhancing detection robustness. However, as shown in Table I and Table II, the APs for pedestrians in WOD and for moderate-level cars and bicycles in KITTI still fall short of the corresponding APs achieved with raw point clouds. This may be because RPKD focuses on transferring detection knowledge via DKD without performing explicit geometric enhancement on lossy compressed point clouds. In future work, we plan to explore geometric enhancement modules at the receiver to recover missing spatial details, thereby further improving the robustness of 3D object detection against non-reflectance compressed point clouds.
@@ V Conclusion
We have presented a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds. During detection training, the designed RCM module assigns the average reflectance of raw point-cloud voxels to corresponding compressed points based on their spatial relationship, providing essential prior labels for reflectance reconstruction. Treating non-reflectance compressed point clouds as weak students, the geometry-based RP module plays a critical role in precise reflectance prediction and object detection within compressed point-cloud detectors. The raw point-cloud detector, with the same structure, transfers the accurate reflectance prediction and object detection knowledge through RKD and DKD, significantly enhancing the robustness of compressed point-cloud detectors. Experimental results on the KITTI dataset and WOD demonstrate that our RPKD framework consistently improves detection performance for compressed point clouds across various code rates and backbone networks.

### v2

@@ IV-B Experimental Results
Table I presents the performance improvement of our RPKD framework on compressed point clouds, using the PV-RCNN, Voxel-RCNN, and SECOND backbones on the KITTI validation split. For PCC-S-C12, PCC-S-C11, and PCC-S-C10 compressed point clouds, RPKD-PV increases mAP by 3.33, 4.23, and 3.65, respectively; RPKD-Voxel improves mAP by 2.72, 2.98, and 2.97; and RPKD-SECOND achieves mAP gains of 2.62, 3.28, and 4.8. As a result, it reveals that our RPKD framework’s detectors exhibit consistent effectiveness and robustness across various code rates and backbones. Additionally, for compressed point clouds with the same code-rate order, RPKD-PV increases the AP values for moderate-level pedestrians by 9.02, 11.98, and 9.66, respectively; RPKD-Voxel boosts the AP values by 8.12, 8.06, and 7.87; and RPKD-SECOND improves the AP values by 3.39, 5.22, and 7.05. These findings highlight the remarkable enhancement in 3D detector robustness for challenging small-sized pedestrians in compressed point clouds.
@@
As depicted in Fig. 6(a), we compared the detection performance of PV-RCNN (STS-C), SparseKD-PV (CDTS), and RPKD-PV (CDTS) on the KITTI compressed data across multiple code rates, using the PV-RCNN backbone. RPKD-PV outperforms SparseKD-PV at all code rates, achieving the highest mAP values. Fig. 6(b) shows the generalization performance of our RPKD framework on the Voxel-RCNN and SECOND backbones. Both RPKD-Voxel (CDTS) and RPKD-SECOND (CDTS) surpass their respective baselines across all code rates, with the two-stage RPKD-Voxel achieving higher mAP than the single-stage RPKD-SECOND. These results vividly illustrate the exceptional robustness and significant performance improvements of our method for 3D object detection on compressed point clouds.
To assess the generalization ability of our RPKD framework, we evaluated its compressed data detection performance with the PV-RCNN backbone on DAIR-V2X-V. Table III further shows that RPKD-PV increases mAP by 4.77, 3.34, and 1.28 on the DAIR-V2X-V validation split. These experimental results demonstrate that our RPKD framework consistently delivers strong performance across varying code rates, backbone networks, and detection datasets.
@@
| Voxel-wise | Point-wise | w/o RP | w/ RP | |||
@@ IV-C2 Critical Elements of Our RPKD
As shown in Table IV, we evaluated the impact of critical elements in our RPKD framework on the compressed data detection performance. In the first two rows, we compared the pretraining performance of compressed point-cloud detectors, where refers to the voxel-based averaging operation in the RCM module. denotes the process of screening foreground compressed points for RKD. In the third row, we used the matched raw point reflectance as the distillation label during point-wise RKD, without performing the voxel-based . In contrast, the fourth row employed the matched reflectance of raw point-cloud voxels as the distillation label during voxel-wise RKD, with the voxel-based . The sixth and final rows assessed the effect of incorporating the RP mechanism into the teacher model, which provides the first-stage proposals for DKD. As indicated in the last row, our RPKD framework achieves the best performance under these settings.
@@ IV-C3 Designed Modules of Our RPKD
As shown in Table V, we evaluated the designed module’s effectiveness for 3D object detection on compressed point clouds. The first row reports the baseline performance of PV-RCNN (STS-C) on PCC-S-C11. We then incrementally integrated the RP module, RKD and DKD into the detector and recorded the corresponding mAP values. Each module contributes to performance improvement, with the highest mAP achieved when all modules are combined.
| Method | Stage | Car 3D | Ped. 3D | Cyc. 3D | Overall |
| PointPillars (STS-C) [27] | One | 80.45 | 45.22 | 64.76 | 63.48 |
| CenterPoint (STS-C) [28] | 80.87 | 50.74 | 65.88 | 65.83 | |
| SECOND (STS-C) [25] | 83.06 | 47.80 | 66.81 | 65.89 | |
| RPKD-SECOND (CDTS) | 83.72 | 52.89 | 70.90 | 69.17 | |
| PV-RCNN (STS-C) [23] | Two | 84.48 | 51.61 | 72.03 | 69.37 |
| PV-RCNN++ (STS-C) [26] | 84.56 | 58.88 | 71.88 | 71.77 | |
| Voxel-RCNN (STS-C) [24] | 84.97 | 55.19 | 73.41 | 71.19 | |
| RPKD-PV (CDTS) | 84.76 | 63.38 | 72.66 | 73.60 | |
| RPKD-Voxel (CDTS) | 84.95 | 62.86 | 74.70 | 74.17 |
IV-C4 Comparison of Single-Stage and Two-Stage Methods
In Table VI, we compared our method against advanced single-stage detectors (PointPillars, CenterPoint, and SECOND) and two-stage detectors (PV-RCNN, PV-RCNN++, and Voxel-RCNN) on PCC-S-C11. Our RPKD (CDTS) improves overall mAP by 3.28, 4.23, and 2.98 on the SECOND, PV-RCNN, and Voxel-RCNN backbones, respectively. Notably, RPKD-Voxel (CDTS) achieves the highest overall mAP across all detection methods. These results demonstrate that our method enhances detection performance for both single-stage and two-stage detectors, with two-stage models exhibiting stronger detection robustness.
IV-C5 Visual Results on DAIR-V2X-V
Fig. 7 presents a visual comparison on the DAIR-V2X-V validation split. RPKD-PV (CDTS) produces fewer false positives and missed detections across multiple code rates, confirming its robustness. As shown in Fig. 8, we used the maximum reflectance value from either raw or compressed data to color the corresponding visual points within each BEV grid. For PCC-S-C12 and PCC-S-C11, the predicted reflectance from compressed data closely matches the actual reflectance from raw data in the primary reflective regions of foreground objects, highlighting the effectiveness of the RP module. However, for PCC-S-C10, the limited number and low density of compressed points lead to a reduction in accuracy, diminishing the reliability of the reflectance prediction.
IV-D Analysis of Compression Transmission
In large-scale V2X scenarios, early fusion is constrained by the high bandwidth required to transmit raw LiDAR data. To address this, we employed voxel-based lossy geometric compression and omitted reflectance encoding, thereby maximizing bandwidth savings. For compression transmission–based cooperative perception, our analysis and evaluation proceed as follows.
IV-D1 Memory and Bandwidth of Raw LiDAR Data
Onboard LiDAR sensors capture point cloud data consisting of single-precision floating-point coordinates (X, Y, Z) and 8-bit unsigned integer reflectance values. The memory usage of a coordinate and reflectance component is denoted as . Accordingly, the memory per raw point is . For LiDAR sensors in the KITTI and DAIR-V2X-V datasets, the number of points acquired per second is . The corresponding transmission bandwidth is , with the detailed requirements summarized in Table VII.
| Dataset | Name | Symbol | Value | Unit |
| Common Parameters | Single Coordinate Memory | 32 | Bit | |
| Reflectance Memory | 8 | Bit | ||
| Single Point Memory | 104 | Bit | ||
| KITTI | Points Per Second | - | ||
| Transmission Bandwidth | 135.2 | Mbps | ||
| DAIR-V2X-V | Points Per Second | - | ||
| Transmission Bandwidth | 74.88 | Mbps |
In vehicular networking, prior studies [59, 60, 61] report that the channel capacity for connected vehicles typically ranges from 30 Mbps to 1 Gbps under varying traffic conditions. Within these limits, transmitting raw point clouds across multiple vehicles overloads both uplink and downlink channels, making real-time data sharing impractical. This highlights the necessity of point cloud compression, where data are compressed and encoded before transmission to enable efficient and feasible cooperative perception.
IV-D2 Bandwidth Calculation for Compressed Point-Cloud Transmission
Existing LiDAR point cloud compression modes can be classified into three categories based on the degree of information loss: lossless compression, lossy compression, and lossy geometry compression. In G-PCC [5], the number of points per second in losslessly compressed point clouds remains identical to that of the raw point clouds, denoted as . The corresponding transmission bandwidth is defined as:
where and denote the code rates for coordinates and reflectance in losslessly compressed point clouds, respectively. In G-PCC–based lossy compression, the number of compressed points decreases according to the octree depth. Let denote the number of points per second in lossy compressed point clouds, thus the transmission bandwidth is expressed as:
where and represent the code rates for coordinates and reflectance in lossy compressed point clouds, respectively. Moreover, lossy geometry compression further reduces transmission bandwidth by discarding reflectance encoding. Assuming the number of points per second remains , the transmission bandwidth for lossy geometry compressed point clouds is obtained by the following formula:
IV-D3 Evaluation for Multi-Vehicle Communication
Based on G-PCC, Table VIII reports the multi-vehicle communication performance of different compression modes on the KITTI dataset. The key metrics include coordinate Bpp, reflectance Bpp, transmission bandwidth, and channel margin. Since coordinate and reflectance code rates lack a fixed proportional relationship, we selected representative values from prior compression studies [9, 7] under medium compression ratio settings, corresponding to a point retention rate of 0.85. The channel capacity for connected vehicles was set to 200 Mbps, a level achievable in most traffic scenarios. The channel margin for a given vehicle is defined as:
where and denote the bandwidth required for receiving and sending point clouds with a single neighboring vehicle, and is the number of connected vehicles, all adopting the same compression setting.
| Compression Mode | Point Retention Rate | Points Per Second | Coordinate Bpp | Reflectance Bpp | Bandwidth (Mbps) | Channel Capacity (Mbps) | Channel Margin (Mbps) | |
| Uncompressed | 1.0 | - | - | 135.2 | 200 | -70.4 | -2504.0 | |
| Lossless | 1.0 | 20.22 | 4.76 | 32.47 | 200 | 135.06 | -449.4 | |
| Lossy | 0.85 | 3.81 | 1.68 | 6.04 | 200 | 187.92 | 79.2 | |
| Lossy Geometry | 0.85 | 3.81 | - | 4.19 | 200 | 191.62 | 116.2 | |
As shown in Table VIII, lossless, lossy, and lossy geometry compression reduce transmission bandwidth by factors of 3, 21, and 31, respectively. For two-vehicle communication (), the corresponding channel margin ratios are 67.5%, 94.0%, and 95.8%. These results confirm that point cloud compression effectively alleviates bandwidth demands, with lossy geometry compression achieving the lowest bandwidth and highest margin. When scaled to larger communication scenarios (), the channel margin ratios for lossy and lossy geometry compression remain at 39.6% and 58.1%, respectively, underscoring the advantage of discarding reflectance encoding in large-scale cooperative perception.
IV-D4 Evaluation for Compression Transmission on DAIR-V2X-V
In real V2X scenarios, we assessed the effectiveness of lossy geometry compression and transmission on the DAIR-V2X-V dataset, using metrics such as compression ratio, coordinate Bpp, and transmission bandwidth. The LiDAR sensor in this dataset operates with 40 beams at a 10 Hz capture frequency, producing approximately points per second. As reported in Table IX, PCC-S-C12, PCC-S-C11, and PCC-S-C10 achieve compression ratios of 24.55, 45.39, and 98.74, respectively, reducing transmission bandwidth by factors of 34, 94, and 373. These findings demonstrate that the PCC-S lossy geometry compression substantially reduces transmission costs in practical V2X environments, enabling real-time sharing of LiDAR data.
| Data Type | Points Per Second | Compression Ratio | Coordinate Bpp | Bandwidth (Mbps) |
| Raw Data | - | - | 74.88 | |
| PCC-S-C12 | 24.55 | 4.01 | 2.17 | |
| PCC-S-C11 | 45.39 | 2.13 | 0.79 | |
| PCC-S-C10 | 98.74 | 0.96 | 0.2 |
IV-E Discussion of Limitation and Future Work
For non-reflectance compressed data, we supplemented the missing reflectance information using the RP module and RKD, thereby enhancing detection robustness. Nevertheless, as shown in Table I, the APs for moderate-level cars and bicycles in KITTI still lag behind those achieved on raw data. A likely reason is that our RPKD framework emphasizes transferring detection knowledge via DKD without explicitly enhancing the geometry of lossy compressed point clouds.
Moreover, this study represents an initial step toward compression transmission-based collaborative perception, with the benchmark task limited to single-view 3D object detection. In real-world collaborative scenarios, connected vehicles will need to integrate locally collected raw data with decoded compressed data to expand perception range and improve detection accuracy. Therefore, effectively leveraging heterogeneous data of varying quality for registration, fusion, and 3D object detection remains a critical open challenge.
In future work, we plan to develop geometric enhancement modules at the receiver to recover missing spatial details, further strengthening the robustness of 3D object detection on non-reflectance compressed data. Beyond this, we aim to advance collaborative 3D object detection by jointly exploiting raw and enhanced compressed data, thereby pushing forward the development of compression transmission-based collaborative perception.
@@ V Conclusion
We have presented a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds. During training, the RCM module assigns the average reflectance of raw point-cloud voxels to corresponding compressed points based on their spatial relationships, providing essential prior labels for reflectance reconstruction. By treating non-reflectance compressed point clouds as weak data, the geometry-based RP module plays a critical role in accurate reflectance prediction and object detection within compressed point-cloud detectors. Building on CDTS, the raw point-cloud detector with the same structure transfers distillation knowledge through RKD and DKD, significantly boosting the robustness of compressed point-cloud detectors. Experimental results on the KITTI and DAIR-V2X-V datasets demonstrate that our RPKD framework consistently improves detection performance across various code rates and backbone networks.